OpenStudy (anonymous):

hello guyz i need help 4a^2 + 4a -3 = 0 2z^2 - z - 6 = 0 solve by completing the square

OpenStudy (owlcoffee):

phew, It's been a long time since I didin't complete a square, but let's analyze it. by definition: \[a ^{2}+2ab+b ^{2}=(a+b)^{2}\] But we'll center in this little chunk: \[(a+b)^{2}=0\] So by hankelian property we know: \[a+b=0 =>a=-b\] So, not so clear, but I'll say, this little fellow must have roots, that respect the general formula. Meaning that, I could use the expanded form to find those roots and these should be double. But, not so long ago, there was a little trick for these, I'll use the first one to show it to you. \[4a ^{2}+4a-3=0\] I'll move the 3 to the other side of the equality, and then divide both sides by 4: \[a ^{2}+a=\frac{ 3 }{ 4 }\] Now, I shall take the simple exponent a, and divide it by 2, and add that amount to both sides: \[a ^{2}+a+\frac{ 1 }{ 4 }=\frac{ 3 }{ 4 }+\frac{ 1 }{ 4 }\] Now I have a completed square on one side, and a fractional addition on the other, let's solve them: \[(a+\frac{ 1 }{ 2 })^{2}=1\] Now, let's take square root, and knowing that square root of 1, is 1: \[a+\frac{ 1 }{ 2 }=1\] \[a=1-\frac{ 1 }{ 2 }\] \[a=\frac{ 1 }{ 2 }\]

OpenStudy (owlcoffee):

sorry, forgot the: \[a=\pm \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

hey @owlcoffee im stuck here can you help me \[z ^{2}-\frac{ z }{ 2 }=3\]

OpenStudy (anonymous):

@owlcoffee thanks by the way on explaining how to complete the square

OpenStudy (anonymous):

@Abdulhameed any ideas?

OpenStudy (anonymous):

\[x ^{2}-\frac{ x }{ 2}=3\] \[\frac{ 2x ^{2} -x}{2 }=3\] \[6=2x ^{2}-x\] \[2x ^{2}-x-6=0\] \[(x-2)(2x+3)=0\] either x-2=0 then x=2 or 2x+3=0 then 2x=-3 , x=-3/2

OpenStudy (anonymous):

@denizen

OpenStudy (owlcoffee):

\[x ^{2}-\frac{ x }{ 2 }=3\] Same method, since x^2 already has a coefficent of 1, we'll divide the x term by 2, square it, and add it on both sides: \[x ^{2}-\frac{ 1 }{ 2 }x=3\] \[x ^{2}-\frac{ 1 }{ 2 }x+\frac{ 1 }{ 16 }=3+\frac{ 1 }{ 16 }\] Completing the squares and doing the fractionary addition on the right side: \[(x-\frac{ 1 }{ 4 })^{2}=\frac{ 49 }{ 16 }\] Now let's solve for x: \[x-\frac{ 1 }{ 4 }=(+/-)\sqrt{\frac{ 49 }{ 16 }}\] I will not prove it, (because I'm sleepy, haven't slept in like... 2 days) but the square root of a fraction is equal to the square root of numerator and denominator. Square root of 49 is 7, and square root of 16 is 4. so: \[x=\frac{ 1 }{ 4 }\pm \frac{ 7 }{ 4 }\] or, for more confort: \[x=\frac{ 1\pm7 }{ 4 }\] Please, recheck what I did, because I have a lack of sleep and when I'm sleepy I make huge mistakes like 4/2=1/2 (actually did that on a test, I'm such a fail ): )